OpenStudy (curry):

find the angle between the 2 vectors u = <5cos30,5sin30> v=

OpenStudy (anonymous):

It should just end up being the same as the angle between 30 and 240.

OpenStudy (curry):

i know how to solve it , cause i implified to arccos=(rad3)/10 but would u use 150 or 210

OpenStudy (curry):

i thought itt was 210 tgoo but the book syas 150

OpenStudy (anonymous):

Oh, I get the problem. When it says the angle between the two vectors, it wants the smallest angle. Look at this picture:

OpenStudy (anonymous):

\[\cos \theta= <u,v> \div ||v|| ||u||\]

OpenStudy (anonymous):